OpenStudy (anonymous):

Find the constant sum for an ellipse with foci F1 (2, 0), F2 (-6, 0) and the point on the ellipse (2, 6).

OpenStudy (anonymous):

I set up the problem like this: √(2-2)^2+(0-6)^2+ √(-6-2)^2+(0-6)^2

OpenStudy (accessdenied):

Looks good to me so far. So then you simplify?

OpenStudy (anonymous):

I simplified it to: √0^2+-6^2+√-8^2+-6^2 √0+36+√64+36 √36+100 =√136 11.66

OpenStudy (accessdenied):

I agree with the first few steps up until \( \sqrt{0 + 36} + \sqrt{64 + 36} = \sqrt{36 + 100} \) Square roots do not combine by addition. They are essentially treated like exponents and can only combine when multiplied. \(\sqrt{a}\sqrt{b} = \sqrt{ab} \)

OpenStudy (accessdenied):

Instead, you have \( \sqrt{36} + \sqrt{100} \) which can be simplified right there because those are perfect squares underneath square roots.

OpenStudy (anonymous):

Ok so it would be: √36=6 + 10=√100 answer is 16?

OpenStudy (accessdenied):

That looks better to me, yes. :)

OpenStudy (anonymous):

Thank you :) I appreciate the help

OpenStudy (accessdenied):

you're welcome!